The length of the rectangle is 1 inches more then the width. The area is 30 square inches. Find the dimensions.

Answer by checkley79(3341)

A=LW
L=W+1
30=(W+1)W
30=W^2+W
W^2+W-30=0
(W+6)(W-5)=0
W-5=0
W=5 ANS. FOR THE WIDTH.
L=5+1=6 ANS. FOR THE LENGTH.
PROOF
30=6*5
30=30